OpenStudy (anonymous):

A circular cylinder is circumscribed about a right prism having a square base one meter in an edge. The volume of the cylinder is 6.283 cu. m. Find its altitude in meters.

OpenStudy (anonymous):

Use the diagonal of the square to get the radius of the circle and try from there.

OpenStudy (anonymous):

so the diagonal of the square is = the the radius correct me if im wrong d = s x \[\sqrt{2}\] d = 1 x \[\sqrt{2}\] d = \[\sqrt{2}\] ?

OpenStudy (anonymous):

i think im wrong. :)) no. im wrong. haha

OpenStudy (anonymous):

Yeah, I mistyped. You have the diameter of the cylinder. Sorry, don't let that mess you up. Now that you have the diameter and the volume, use the formula to find the height. No, you are correct. a^2 + b^2 = c^2 proves you.

OpenStudy (anonymous):

I'm kinda confused. What do you mean by the diameter of the cylinder?

jhonyy9 (jhonyy9):

hi ! i think that for more easy understanding you need drawing it

OpenStudy (anonymous):

I got it! Thanks :D

OpenStudy (anonymous):

So the answer is 4m. :)